# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 7.2. Specification of the Cauchy data
The Cauchy data $(g,k)$ on $\widehat{\mathcal{M}}_{t=-a}$, consisting of the induced Riemannian metric $g$ and second fundamental form $k$, are specified as follows:
@@
    Interior data. On $\mathcal{M}_{1}$ we prescribe smooth $C^{\infty}$ data $(g,k)$ satisfying
    (195) $\operatorname{Rad}(\mathcal{M}_{1})<\frac{3\pi}{2\,c_{\partial\mathcal{M_{1}}}},$
    where $c_{\partial\mathcal{M}_{1}}$ denotes the generalized mean curvature coefficient of $\partial\mathcal{M}_{1}$ in the sense of Schoen–Yau. Existence of such data follows from the quasi-local analysis of [38, 37].
@@
    Matching data in the characteristic region. On $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$, the data are induced from the characteristic development on the double-null slab $\mathcal{D}_{-a,\epsilon}$. Well-posedness of the characteristic problem on this region was established in the previous section.
@@
    Asymptotically Kerr exterior. On $\mathcal{M}_{3}$, we prescribe smooth data $(g,k)$ asymptotic to a Kerr slice with prescribed mass $m>0$ and angular momentum $J$. This ensures asymptotic flatness.
A gluing construction in the spirit of [20, 23, 10] yields a smooth, asymptotically flat vacuum Cauchy pair $(g,k)$ on $\widehat{\mathcal{M}}_{t=-a}$. Complete geometric–analytic details of this construction will appear in [28].
@@ 7.3. Setup, decomposition, constraints, and gauges
Let (^,g^)($\widehat{\mathcal{M}}$$\widehat${g}) be the semi-globally constructed Lorentzian manifold equipped with a canonical double–null foliation $(u,\underline{u})$ and Cauchy time $t=u+\underline{u}$. Fix $a\gg 1$ and a small $0<\epsilon\ll 1$. On the initial slice $\widehat{\mathcal{M}}_{t=-a}$ set
$\widehat{\mathcal{M}}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
with $\mathcal{M}_{1}=\widehat{\mathcal{M}}_{\mathrm{int}}$, $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$ (the double–null slab), and $\mathcal{M}_{3}=\widehat{\mathcal{M}}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2})$. We must solve on $\widehat{\mathcal{M}}_{t=-a}$ the vacuum constraint system
@@ 7.4. Interior data on $\mathcal{M}_{1}$: strict Yau–radius subcriticality with boundary control
Let $\Omega:=\mathcal{M}_{1}$ and denote its outer boundary by $\partial\Omega=\partial\widehat{\mathcal{M}}_{\mathrm{int}}$. Let $H$ be the mean curvature of $\partial\Omega$ in the initial slice and $p\equiv k$ the slice’s second fundamental form in spacetime. Fix a strictly positive boundary gap
(197) $c_{\ast}:=\inf_{\partial\Omega}\bigl(H-|\mbox{tr}_{\partial\Omega}k|\bigr)>0,$
and choose a target radius bound $0<\text{Rad}(\Omega)\leq R_{\ast}<R_{\mathrm{crit}}(c_{\ast})$ strictly below Yau’s critical threshold (quantified below).
Boundary barrier via Yau’s brane functional.
Let $f>0$ be the first eigenfunction for the mixed operator $\Delta+\tfrac{1}{2}R-h$ with Robin data $\partial_{\nu}f+kf=0$ on $\partial\Omega$ (see [38] for details). For $\Gamma\subset\partial\Omega$ realizing $\text{Rad}(\Omega)$ up to $o(1)$, and a spanning disk $\Sigma\subset\Omega$ with $\partial\Sigma=\Gamma$, set
$\mathcal{L}_{f}(\Sigma)=\int_{\Sigma}f-c\int_{\Omega_{\Sigma}}f,\qquad c\in(0,c_{\ast}].$
[…]
$\min_{\partial\Omega}\bigl(H(\tilde{g})-|\mbox{tr}_{\partial\Omega}\tilde{k}|\bigr)\geq c_{\ast}>0,\qquad\text{Rad}(\Omega)\leq R_{\ast}<R_{\mathrm{crit}}(c_{\ast}),$
so $\mathcal{M}_{1}$ contains no apparent horizon on t=at=-a. Moreover, these inequalities are stable under the smoothing that we will do across the interfaces.
7.5. Characteristic matching on $\mathcal{M}_{2}$: importing the double–null data
On $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$, take the data induced from the double–null development on the slab $\mathcal{D}_{-a,\epsilon}$, in the normalized frame
@@ 7.5. Characteristic matching on $\mathcal{M}_{2}$: importing the double–null data
$\|\widehat{\chi}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim|u|^{-1},\qquad\|\widehat{\underline{\chi}}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{1/2}|u|^{-2},$
$|\mbox{tr}\chi|\lesssim|u|^{-1},\quad|\mbox{tr}\underline{\chi}|\lesssim|u|^{-1},\quad\|\eta\|_{L^{\infty}}+\|\underline{\eta}\|_{L^{\infty}}\lesssim a^{1/2}|u|^{-2},\quad|\omega|+|\underline{\omega}|\lesssim|u|^{-1}.$
These are propagated from the double–null hierarchy and are compatible with the transport system
$\nabla_{3}\mbox{tr}\chi+\tfrac{1}{2}(\mbox{tr}\chi)^{2}=-|\widehat{\chi}|^{2}-2\underline{\omega}\,\mbox{tr}\chi,\qquad\nabla_{4}\mbox{tr}\underline{\chi}+\tfrac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-|\widehat{\underline{\chi}}|^{2}-2\omega\,\mbox{tr}\underline{\chi},$
together with the $\nabla_{4}$$\nabla_{3}$ equations for $\widehat{\chi}$ and $\widehat{\underline{\chi}}$. In particular, on the interface sphere $S_{-a,\underline{u}}$ the combination
@@
7.6. Asymptotically Kerr exterior on $\mathcal{M}_{3}$ and compactly supported correction
Choose a large coordinate sphere that encloses $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ and fix asymptotically flat coordinates outside. Pick a member of the admissible Kerr family (mass $m>0$, angular momentum $J$ and Poincaré parameters) and carry out the Corvino–Schoen gluing: smoothly splice $(g,k)$ to the chosen Kerr slice in an annulus $A_{R}=\{R<r<2R\}$, then perturb by a compactly supported deformation inside $A_{R}$ to annihilate the ten–dimensional cokernel of the linearized constraint operator in $W^{2,p}_{-\delta}\times W^{1,p}_{-1-\delta}$. The surjectivity of the linearization in these scales and the degree–argument for the parameter map $\Theta$ produce a solution $(g^{\sharp},\pi^{\sharp})$ that equals Kerr outside $|x|\geq 2R$ and equals the original data inside $|x|\leq R$, while solving (196) globally.
7.7. Smoothing across interfaces and final properties on t=at=-a
Using a collar–supported partition of unity on the 2-spheres separating $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$, we ensure $C^{\infty}$ matching of the slice’s first and second fundamental forms, and re-solve the constraints by compactly supported corrections that do not alter $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ nor the exterior Kerr end. The resulting global Cauchy data $(g,k)$ on $\widehat{\mathcal{M}}_{t=-a}$ satisfy:
  1. (1)
    $(g,k)$ are smooth and vacuum everywhere and asymptotically Kerr outside a compact set;
  2. (2)
    $\mathcal{M}_{1}$ is Yau–radius subcritical with boundary gap $c_{\ast}>0$; hence no MOTS occur on t=at=-a;
  3. (3)
    the characteristic shear content on $\mathcal{M}_{2}$ ($\widehat{\underline{\chi}}$) is large in the prescribed sense; the future slice $t=-a+\epsilon$ then satisfies a strengthened boundary–curvature inequality $c_{S_{-a,\epsilon}}>\tfrac{3\pi}{2\text{Rad}(\cdot)}$ for the interior domain, forcing an apparent horizon in its interior in the future development.
7.8. Proof notes for the boundary barrier and the Yau radius threshold
Let $f$ solve the mixed eigenvalue problem $\partial_{\nu}f+kf=0$ on $\partial\Omega$ as stated in [38]. For $\Gamma\subset\partial\Omega$ realizing $\text{Rad}(\Omega)$, minimizing $\mathcal{L}_{f}(\Sigma)$ among disks $\Sigma$ with $\partial\Sigma=\Gamma$ yields a stationary surface obeying a mean–curvature equation whose first and second variations control the concentrating competitors inside the tubular neighborhood of $\partial\Omega$. Using the barrier built from the distance function $d(x,\partial\Omega)$ and profiles $\varphi(d)$ with $\varphi^{\prime}(\varepsilon)=\infty$, one enforces that minimizers stay away from $\partial\Omega$, and the second variation produces the quantitative inequality linking $c_{\ast}$ and $\text{Rad}(\Omega)$. Choosing $R_{\ast}<R_{\mathrm{crit}}(c_{\ast})$ excludes MOTS on t=at=-a. Now note the following vital points that are made explicit.
  • The interior construction produces $(g,k)$ with a fixed positive gap $H-|\mbox{tr}_{\partial\Omega}k|\geq c_{\ast}$ and with $\text{Rad}(\Omega)$ prescribed strictly below the critical Yau threshold.
@@ 7.8. Proof notes for the boundary barrier and the Yau radius threshold
  • The exterior gluing gives an exact Kerr outside a compact set without spoiling the interior Yau gap or the characteristic content.
Now we briefly discuss the constraints. ∎
7.9. Constraint map and local deformation
Let $\Phi(g,\pi)=\big(\text{Scal}(g)+(\mbox{tr}_{g}\pi)^{2}-|\pi|_{g}^{2},\ \text{Div}_{g}\pi-d(\mbox{tr}_{g}\pi)\big)$ denote the vacuum constraint operator on a 33–manifold. Its linearization $D\Phi_{(g,\pi)}$ at an asymptotically flat (AF) seed and the formal $L^{2}$–adjoint $D\Phi_{(g,\pi)}^{*}$ are overdetermined–elliptic. On weighted Sobolev scales $W^{2,p}_{-\delta}\times W^{1,p}_{-1-\delta}$ ($p>3$, $0<\delta<1$), the adjoint has surjective range modulo a finite–dimensional obstruction space [25, 26]
$\mathcal{K}\cong\mathbb{R}^{10},$
canonically identified with Killing initial data corresponding to the Poincaré charges $(E,P,J,C)$ at infinity. Compactly supported solutions of
$D\Phi^{*}_{(g,\pi)}(u,X)=F\quad\text{on a domain }U\Subset M$
exist and yield compactly supported corrections to $(g,\pi)$ after a quadratic iteration, provided the $\mathcal{K}$–projection of the residual is killed by parameter variation in an admissible model family (Kerr) on the outer end.
7.10. Regionwise seeds
Fix times $t_{1}=-a$, $t_{2}=-a+\epsilon$, $\epsilon>0$ small. Decompose the initial slice
(198) $\widehat{\mathcal{M}}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
with interfaces along smooth 22–spheres.
(I) Interior seed on $\mathcal{M}_{1}$ with prescribed Yau radius as discussed previously.
Let $\Omega:=\mathcal{M}_{1}$ with outer boundary $\partial\Omega$. Choose a smooth Riemannian pair $(g^{\rm in},k^{\rm in})$ on $\Omega$ such that:
  1. (i)
    Boundary gap & Yau radius. On $\partial\Omega$,

### v4

@@ 6.2. Specification of the Cauchy data
The Cauchy data $(g,k)$ on $\mathcal{M}_{t=-a}$, consisting of the induced Riemannian metric $g$ and second fundamental form $k$, are specified as follows:
@@
    Interior data.
@@
    Matching data in the characteristic region. On $\mathcal{M}_{2}=\mathcal{M}_{t=-a}\cap D_{a,1}$, the data are induced from the characteristic development on the double-null slab $D_{a,1}$. Well-posedness of the characteristic problem on this region was established in the previous section.
@@
    Asymptotically Kerr exterior. On $\mathcal{M}_{3}$, we prescribe smooth data $(g,k)$ asymptotic to a Kerr slice with prescribed mass $m=O(a^{1/2})$ and angular momentum $J=a$. This ensures asymptotic flatness.
A gluing construction in the spirit of [22, 28, 11] yields a smooth, asymptotically flat vacuum Cauchy pair $(g,k)$ on $\mathcal{M}_{t=-a}$.
@@ 6.3. Setup, decomposition, constraints, and gauges
Let (,g^)($\mathcal{M}$$\widehat${g}) be the semi-globally constructed Lorentzian manifold equipped with a canonical double–null foliation $(u,\underline{u})$ and Cauchy time $t=u+\underline{u}$. Fix $a\gg 1$. On the initial slice $\mathcal{M}_{t=-a}$ set
$\mathcal{M}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
with $\mathcal{M}_{1}=\mathcal{M}_{1}$, $\mathcal{M}_{2}=\mathcal{M}_{t=-a}\cap\mathcal{D}_{-a,1}$ (the double–null slab), and $\mathcal{M}_{3}=\mathcal{M}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2})$. We must solve on $\mathcal{M}_{t=-a}$ the vacuum constraint system
@@
On $\mathcal{M}_{2}=\mathcal{M}_{t=-a}\cap D_{a,1}$, take the data induced from the double–null development on the slab $\mathcal{D}_{-a,\epsilon}$, in the normalized frame
@@
$\|\widehat{\chi}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}|u|^{-1},\qquad\|\widehat{\underline{\chi}}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{\frac{1}{2}}|u|^{-2},$
$|\mbox{tr}\chi|\lesssim|u|^{-1},\quad|\mbox{tr}\underline{\chi}|\lesssim|u|^{-1},\quad\|\eta\|_{L^{\infty}}+\|\underline{\eta}\|_{L^{\infty}}\lesssim a^{\frac{1}{2}}|u|^{-2},\quad|\omega|+|\underline{\omega}|\lesssim a^{\frac{1}{2}}|u|^{-2}.$
In particular, on the interface sphere $S_{-a,\underline{u}}$ the combination
@@
In this section, we complete the proof of Theorem 2.2 by establishing the formation of an MOTS strictly to the future of the Cauchy hypersurface
$\mathcal{M}_{t=-a}\subset\mathcal{M}.$
We work in the spacetime harmonic gauge and solve the vacuum Einstein equations on the slab
$a,-a-1/a+\epsilon]\times\mathcal{M},\qquad 0<\epsilon\ll 1,$
with initial data assembled by a gluing procedure. We then propagate the quasi-local boundary geometry forward in time, ultimately showing that the Schoen–Yau curvature radius condition is violated in the future domain, forcing the existence of an MOTS. In this section, we provide a sketch of the construction of the data.
This is achieved by a refined geometric decomposition and constraint–compatible gluing of the interior Cauchy data. We decompose
$\mathcal{M}_{1}=\widetilde{\mathcal{M}}_{1}\cup\Big(\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}\Big),$
where $\widetilde{\mathcal{M}}_{1}\Subset\mathcal{M}_{1}$ is a compact subdomain with smooth boundary and the complement $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ is a collar (transition) region of uniformly bounded thickness $O(1)$ measured with respect to the background Euclidean metric. The decomposition is chosen so that $\partial\widetilde{\mathcal{M}}_{1}$ lies a fixed positive distance away from the outer interface $\partial\mathcal{M}_{1}$ and from the gluing interface with the characteristic region $D_{a,1}$.
On $\widetilde{\mathcal{M}}_{1}$ we prescribe vacuum constraint data $(g,k)$ satisfying the Einstein constraint equations and the quantitative smallness bounds
(231) $\|k\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-3/2},~\|\partial g\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}$ $\leq Ca^{-1},$
(232) $\|k\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-3/2},~\|\partial g\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}$ $\leq Ca^{-1},$
for some fixed $s\gg 1$, where $H^{s-1}_{\mathrm{ul}}$ denotes the uniformly–local Sobolev space (defined via unit–scale coordinate balls). The same bounds hold for the lapse $N$ and shift $X$,
(233) $\|N-1\|_{L^{\infty}}+\|X\|_{L^{\infty}}+\|N-1\|_{H^{s}_{\mathrm{ul}}}+\|X\|_{H^{s}_{\mathrm{ul}}}\leq Ca^{-3/2},$
so that the data are a quantitatively small perturbation of Euclidean data (in homogeneous Sobolev norm) on each unit ball, with constants independent of the total $H$-radius $O(a)$ of $\mathcal{M}_{1}$. In particular, the local constraint norms remain $O(a^{-3/2})$ while global Sobolev norms may be $O(1)$ due to volume growth; all arguments below are formulated in uniformly–local norms to avoid this scaling loss.
The collar region $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ is used as a gluing zone in which $(g,k)$ is constructed by a constraint–preserving interpolation between the interior data on $\widetilde{\mathcal{M}}_{1}$ and the induced Cauchy data coming from the characteristic development $D_{a,1}$ on $\mathcal{M}_{2}$. Using a Corvino–Schoen type localized deformation together with a partition of unity and solvability of the linearized constraint operator with compact support, one obtains smooth vacuum data on $\mathcal{M}_{1}$ that agree exactly with the interior data on $\widetilde{\mathcal{M}}_{1}$ and with the characteristic–induced data near $\partial\mathcal{M}_{1}$, with all corrections supported strictly inside the collar (a relevant proposition 6.2 and a sketch of the proof. Since the collar has thickness $O(1)$, these localized corrections do not alter any geometric radius quantity of order $a$ except by an additive $O(1)$ error. Consequently, if $\text{Rad}_{g}(\widetilde{\mathcal{M}}_{1})$ denotes the Schoen–Yau (HH–)radius computed with respect to the interior metric $g$, then
(234) $\text{Rad}(\mathcal{M}_{1})=\text{Rad}_{g}(\widetilde{\mathcal{M}}_{1})+O(1),$
with the $O(1)$ constant depending only on the collar geometry and the fixed gluing profile.
Set $\Omega:=\mathcal{M}_{1}$ and denote its outer boundary by $\partial\Omega=\partial\mathcal{M}_{1}$. Let $H$ be the mean curvature of $\partial\Omega$ in the initial slice $(\Omega,g)$ with respect to the outward unit normal, and let $k$ be the second fundamental form of the slice in spacetime. Because the gluing corrections are supported away from $\partial\Omega$, the induced pair $(\gamma,k^{\top})$ on $\partial\Omega$ coincides with that coming from the double–null interface data. In particular, there is a strictly positive barrier gap
(235) $c_{*}:=\inf_{\partial\Omega}\Big(H-|\mbox{tr}_{\partial\Omega}k|\Big)>0.$
Fix a target radius bound
(236) $0<\text{Rad}(\Omega)\leq R_{*}<R_{\mathrm{crit}}(c_{*}),$
strictly below the Schoen–Yau critical radius associated to the gap $c_{*}$.
Following Schoen–Yau, consider the mixed boundary value problem for the operator
$\mathcal{L}:=-\Delta+\tfrac{1}{2}R-h$
on $\Omega$, with Robin boundary condition
$\partial_{\nu}f+(\mbox{tr}_{\partial\Omega}k)\,f=0\quad\text{on }\partial\Omega,$
where $R$ is the scalar curvature of $(\Omega,g)$ and $h$ is the quadratic form in $k$ appearing in the stability operator for marginally outer trapped surfaces (see [44] for the precise formula). Standard elliptic theory yields a strictly positive first eigenfunction $f>0$.
Let $\Gamma\subset\partial\Omega$ be a boundary curve realizing $\text{Rad}(\Omega)$ up to $o(1)$, and let $\Sigma\subset\Omega$ be a spanning disk with $\partial\Sigma=\Gamma$. Define the weighted functional
$\mathcal{L}_{f}(\Sigma)=\int_{\Sigma}f\,d\mu_{\Sigma}-c\int_{\Omega_{\Sigma}}f\,d\mu_{g},\qquad c\in(0,c_{*}],$
where $\Omega_{\Sigma}$ is the region enclosed by $\Sigma$ and $\Gamma$. Using first and second variation formulas for $\mathcal{L}_{f}$, together with the strict boundary barrier (235) and the distance–to–boundary foliation with explicit supersolution/subsolution profiles $\varphi(d)$, one obtains existence of a minimizing surface with $\partial\Omega$ acting as a strict barrier and derives the quantitative curvature–radius inequality of Schoen–Yau type. In particular, if $\text{Rad}(\Omega)<R_{\mathrm{crit}}(c_{*})$, then no marginally outer trapped surface can be contained in $\Omega$.